# To be or not to be, but where?

### v1

@@ Abstract
The identification of physical subsystems in quantum mechanics as compared to classical mechanics poses significant conceptual challenges, especially in the context of quantum gravity. Traditional approaches associate quantum systems with classical ones localized in spacetime, using either Hilbert space factors for finite-dimensional systems or local operator algebras in algebraic quantum field theory. These methods ensure statistical independence for state preparations and measurements. However, canonical linearized quantum gravity disrupts this framework by preventing the formation of gauge-invariant local algebras, thereby undermining the statistical independence required in measurements. This presents a major obstacle for modeling early universe cosmology, gravity-induced-entanglement experiments, and poses a significant roadblock toward a comprehensive theory of quantum gravity. A pivotal shift is proposed: the identification of classical and quantum systems should be dynamically evolving rather than static, opening the possibility of a single-world unitary quantum mechanics. This perspective aligns with the broader aim of understanding how classical spatiotemporal existence emerges from quantum mechanics and connects the measurement problem with quantum gravity.
@@ 1 The mysterious box
One day, you arrive at your lab and find a mysterious box. The box is completely black except for a small display where it says “ENERGY SPECTROMETER”. You try to open the box but cannot quite figure out how. So, you decide to interact with the display, and write down all the possible energy values of whatever is inside the box. As you near the end of this tedious task, all the years spent modeling physical systems give you a hunch that perhaps whatever is inside the box is actually a simple physical system.
@@
You believe there are 4 qubits, each coupled to its neighbor in the $z$-direction, with a magnetic field in the same direction. You compute the energy spectrum of this system, and, to your surprise, there is a pair of $J$ and $h$ for which the spectrum matches exactly with the values displayed on the black box!
Excited, you storm out of the lab and call a friend to see the black box and share your discovery. When you both return to the lab and you are about to show your model, your friend stops you, wanting to figure out the mystery themselves. After some time, they return, having found their own model for what is inside the box, described by the following Hamiltonian:
@@
It is also a model of spin sites! It looks quite similar to your own, except that the coupling between their spin sites is not the same as yours, but what you had as the magnetic field. And their magnetic field is on the $x$-direction, and not over all their spins… also, their first spin is under another magnetic direction on the $z$-direction. How is that possible?
You discuss what is going on trying to understand how you both arrived at different models of what is inside the box that nonetheless matched exactly the energy spectrum. Until your friend has an idea: what if what you both called qubits are somehow related to each other? After some thought, you arrive at the following identification:
@@
which, together with the swapping of the coupling strength and magnetic field, maps one model into the other exactly.
At first, you are not very surprised, as it is quite common in physics to find different models to describe the same phenomenology. However, as you both ponder about this map, you start to become suspicious about what is going on. For instance,
@@
tells you that the spin site 4 in the $\mu$-model in the $z$-direction is somehow mapped into the product of all the spin sites in the $x$-direction in the $\sigma$-model. Meanwhile,
@@
relates the $\mu$-sites in the $x$-direction to pairwise neighbors in the $\sigma$-model in the $z$-direction. The problem is that both you and your friend used the numbers ${1,2,3,4}$ to label spin sites, which are localized in spacetime. However, the spin sites in each description are non-locally related. Each model assumes a different notion of the system, interactions, and spatial distribution of the physical subsystems. But, surely, one model must be the real one, or so you hope.
As you both start to lose patience and commit to discovering what the box hides, you decide to follow the example of particle physicists and grab a hammer to smash it open. Just as you are about to strike the first blow, the box disappears. Nonetheless, the question remains: how can you make sense of what was inside the box?
@@ 2 How space hides in quantum mechanics
The gedanken experiment above illustrates that quantum systems only described by their global Hilbert space, $\mathcal{H}$, a state vector, |$\ket{\Psi}$, and a Hamiltonian, $H$, might not fully characterize its classical properties, such as the notion of a “site” - a location in space. After all, we only had access to the overall energy associated with each of the global states of the system [^4] Once again, we considered a simplified scenario where each energy value displayed in the spectrometer is associated with a non-degenerated energy eigenstate, thus also fixing the size of the Hilbert space. All the eigenvalues together can be used to reconstruct the Hamiltonian matrix in its diagonal form.. This is puzzling because typically, the situation is reversed: we impart classical notions of location into our quantum mechanical descriptions. For example, for every spin site located at position $i$, we assign a $\mathbb{C}^{2}$-Hilbert space factor, such that the global Hilbert space is
$\mathcal{H}=\bigotimes_{i=1}^{n}\mathbb{C}^{2}\,.$ (2.1)
This is just a simple example of a general attitude [^5] We can speculate whether this attitude is intrinsically derived from the quantization program, where classical observables are promoted to quantum operators, and among them, position and momentum take a special role. Had the program been classicalization instead, then the emphasis could have been rather different.: we typically assign a Hilbert space $\mathcal{H}_{\rm system}$ to each classical subsystem in a given spatial (or more generally, spacetime) region [^6] As usual, this story is fairly more complicated, in particular in the context of infinite-dimensional systems and constrained theories, where only the notion of localizability for the observables is guaranteed [1]. We will come back to that in Sec. 5 and 6. $U$. Thus, some notion of space, or at very least, some notion of location in space seems to leak into our quantum-mechanical descriptions. Let’s make that explicit.
@@
Thus, we have two commuting sets of observables. If we also require that they together span all the possible operators associated with the system [^7] Rmember that any operator can be decomposed into a complex sum of two self-adjoint operators, i.e., observables., i.e., $\mathcal{O}^{A}\otimes\mathcal{O}^{B}\cong L(\mathcal{H})$, then it follows [4] that the Hilbert space of the quantum system can be factorized, $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. Thus, our local assignment of Hilbert space factors to parts of a bigger system is born out of macrocausality, operationally implemented in a way that leads to microcausality, and finally leading to the factorization of the Hilbert space.
Now that the physical intuition is properly encoded in the math, let’s be more precise and general. Following [4], we consider a finite-dimensional Hilbert space $\mathcal{H}$ and a collection of subalgebras of observables corresponding to each subsystem satisfying the following properties:
  • Local acessibility: Each $\{\mathcal{O}^{I}\}$ corresponds to a set of controllable observables, where $I$ labels the different systems;
  • Subsystem independence: $[\mathcal{O}^{I},\mathcal{O}^{J}]=0$, where $I\neq J$, which we now understand as a direct consequence of operational macrocausality;
  • Completeness: $\bigvee_{I}\mathcal{O}^{I}\cong L(\mathcal{H})$, there is a minimal inclusion of the subalgebras that is isomorphic to the whole algebra of operators.
@@
By choosing a TPS, we determine how the Hilbert space of the entire quantum mechanical system (like the box in the example above) will be partitioned. However, now we understand that this assignment is intrinsically related to our classical notion of spacetime. On the other hand, if nature is fundamentally quantum mechanical, it might make more sense to start with the global Hilbert space and then find an appropriate decomposition [^8] There are recent works in this direction, for instance [6, 7, 5]. - particularly one that aligns with our classical observables, such as the notion of site 1 is here, site 2 is there. The question, however, is whether there are multiple classical possibilities associated with the same quantum system, as seen in the different site attributions in our gedanken experiment above, and vice-versa.
Once a TPS is introduced, we examine the 3-tuple $(\mathcal{H},H,\ket{\Psi})$ through the composition of each factor $\mathcal{H}_{i}$ (or local subalgebras, $\{\mathcal{O}_{i}\}$). This is exactly what happened in our box experiment, where you and your friend tacitly assumed different TPSs. Each $\mathcal{T}$ considered is not related by local unitaries or permutations of subsystems; instead, the set of variables are non-locally related as shown in (1.3), amounting to a global unitary transformation. Thus, the TPSs are not equivalent. That can also be checked by fixing the factorization of the Hilbert space, $\bigotimes_{i}\mathcal{H}_{i}$ and showing that for each TPSs $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, the operators $T_{1}HT_{1}^{-1}$ and $T_{2}HT_{2}^{-1}$ are not the same up to conjugation by local unitaries, permutations of subsystems, and transposition. Therefore, a change of TPS has an equivalent perspective where the TPS is fixed but the Hamiltonian is unitarily transformed, keeping its spectrum.
@@
You might expect that the issue of identifying the appropriate classical limit, understood as the assignment of Hilbert space factors to spatiotemporally localized systems, arises only when dual descriptions exist. After all, if the box cannot be opened (i.e., we cannot extend its Hamiltonian [8]), we have incompatible classical descriptions of what is in the box. However, the problem runs deeper. Even without dualities, a generic Hilbert space with a generic Hamiltonian will have multiple tensor factorizations, only constrained by the dimension of the Hilbert space. And without further specification, all these decompositions are equally reasonable representations of the quantum-mechanical system.
This quantum democracy has led to recent efforts to identify criteria for singling out a preferred factorization of the Hilbert space [6, 7]. In particular, [5] suggests a criterion based on the fact that for a TPS with $n$ subsystems, whenever $n\gg k$, the set of $k$-local Hamiltonians has measure zero. If one can then prove that dualities are not present, this would imply an almost unique TPS for a $k$-local Hamiltonian. Thus, by considering $k$-locality of the Hamiltonian as an additional structure to $(\mathcal{H},H,\ket{\Psi})$, a preferred factorization is selected. This also determines a preferred way to map quantum subsystems, as factors in Hilbert space, into classical ones, as systems localized in spacetime.
@@ 3 A Tale of Emergence
In this section, we will discuss the concept of emergence, start to clarify the distinction between classical and quantum systems and establish foundations to better understand our mysterious box.
@@
Key to this definition are the terms novel, robust, and new predictions. When we have a description that is novel and robust, in the sense that it stands on its own and provides a way to predict the behavior of physical systems, we typically call it a physical theory. Of course, this also underlies what we mean by a more fundamental description. Therefore, perhaps we should stop here and scrutinize the structure of what we mean by physical theories to make the different flavors of emergence more explicit.
@@ 3.1 The Structure of Physical Theories
  • Underlying state space ($\Omega$): This represents the space of all possible states of a physical system. For example, in Lagrangian mechanics, this is the configuration space, while in Hamiltonian mechanics, it is the phase space. In quantum mechanics, we typically consider the Hilbert space [^9] See footnote 20, though.;
@@
  • Dynamical equation ($U$): Once a state is known for a given degree of freedom, the dynamical equation determines the evolution of the state, defining a trajectory in the state space. In classical mechanics, we have Newton’s, Euler-Lagrange or Hamiltonian equations [^11] In this section, we will assume classical mechanics is governed by Newton’s equations, as this is sufficient to make our point., while in quantum mechanics, we have the Schrödinger or Heisenberg equation.
In the language of [12], these elements are reduced to a triple $T=(S,Q,D)$: the state-space, the quantities, and the dynamics. The reason why we say most physical theories, is because some do not fit this decomposition: e.g., in thermodynamics, the second law does not take the form of a dynamical equation. Nonetheless, to convey our points in the next sections, we will ignore these subtleties.
@@
Now that we have at hands the definition of emergence and the underlying structure of physical theories, we can establish different levels of emergence by considering some examples.
@@ 3.2 Emergence - Level 0: Classical Mechanics and the Center of Mass
$X_{\rm CM}:\left\{\begin{aligned} &\mathds{R}^{3N}\rightarrow\mathds{R}^{3}\\ &\vec{x}_{\rm CM}\equiv\frac{1}{M}\sum_{I=1}^{N}m_{\rm I}\vec{x}_{\rm I}\,,\quad\quad M=\sum^{N}_{I=1}m_{\rm I}\,,\end{aligned}\right.$
@@
Dynamical Equations $\vec{F}(\vec{x}_{\rm I})=m_{\rm I}\frac{d\vec{v}_{\rm I}}{dt}$ $\vec{F}(\vec{x}_{\rm CM})=M\frac{d\vec{v}_{\rm CM}}{dt}$
@@ 3.2.1   Does the center of mass exist?
As we aim to build our intuition around ontology as emergence unfolds, we can begin by asking some fundamental questions. There is plenty of literature around this topic [13, e.g. and references therein], but we just want to highlight a few things to slowly build our intuition for later on.
First, in some situations, the center of mass defines a point that lies within the composite object it represents, making it materially instantiated. In such cases, there is literally a material point with its own position and velocity vector that follows the center-of-mass trajectory. However, this is not always the case. Often, the center of mass does not correspond to any actual material point within the defining object. From a substantivalist perspective, all points in space exist in an absolute sense, so the center of mass trajectory simply labels some of these points.
Rather than focusing on material instantiation or substantivalism to argue for the existence of the center of mass, we can adopt the perspective supported by our definition of emergence. As long as new entities are novel and stable enough to yield reliable predictions, such as the case for the center of mass, we can attribute emergent ontological status to them, similar to Wallace’s implementation [13] of Dennett’s real patterns [14] in physics.

### v2

@@ Abstract
The identification of physical subsystems in quantum mechanics as compared to classical mechanics poses significant conceptual challenges, especially in the context of quantum gravity. Traditional approaches associate quantum systems with classical ones localized in spacetime, using either Hilbert space factors for finite-dimensional systems or local operator algebras in algebraic quantum field theory. These methods ensure statistical independence for state preparations and measurements. However, covariant linearized quantum gravity disrupts this framework by preventing the formation of gauge-invariant local algebras, thereby undermining statistical independence. This presents a major obstacle for modeling early universe cosmology and gravity-induced entanglement experiments, and poses a significant roadblock toward a comprehensive theory of quantum gravity. A pivotal shift is proposed: the identification between classical and quantum systems should be dynamically evolving rather than fixed, opening the possibility of a single-world unitary quantum mechanics. This perspective aligns with the broader aim of understanding how classical spatiotemporal existence emerges from quantum mechanics and connects the measurement problem with quantum gravity.
@@ 1 The mysterious box
One day, you arrive at your lab and find a mysterious box. The box is completely black except for a small display where it says “ENERGY SPECTROMETER”. You try to open the box, but cannot quite figure out how. So, you decide to interact with the display and write down all the possible energy values of whatever is inside the box. As you near the end of this tedious task, all the years spent modeling physical systems give you a hunch that perhaps whatever is inside the box is a simple physical system.
@@
You believe that there are 4 qubits, each coupled to its neighbor in the $z$-direction, with a magnetic field in the same direction. You compute the energy spectrum of this system, and, to your surprise, there is a pair $\{J,h\}$ for which the spectrum matches exactly the values displayed in the black box.
Excited, you leave and call a friend to see the black box and share your discovery. When you both return to the lab and you are about to show your model, your friend stops you, wanting to figure out the mystery themselves. After some time, they return, having found their model for what is inside the box, described by the following Hamiltonian:
@@
It is also a model of spin sites! It looks quite similar to your own, except that the coupling between their spin sites is not the same as yours, but what you had as the magnetic field. And their magnetic field is along the $x$-direction, and not over all their spins… also, their first spin is under another magnetic field in the $z$-direction. How is that possible?
You discuss what is going on, trying to understand how you both arrived at different models of what is inside the box that nonetheless matched exactly the energy spectrum. Until your friend has an idea: what if what you both called qubits are somehow related to each other? After some thought, you arrive at the following identification:
@@
which, together with the swapping of the coupling strength and magnetic field, maps one model exactly onto the other one.
At first, you are not very surprised, as it is quite common in physics to find different models to describe the same phenomenology. However, as you both ponder this map, you start to become suspicious of what is going on. For instance,
@@
tells you that spin site $4$ in the $\mu$-model in the $z$-direction is somehow mapped onto the product of all the spin sites in the $x$-direction in the $\sigma$-model. Meanwhile,
@@
relates the $\mu$-sites in the $x$-direction to pairwise neighbors in the $\sigma$-model in the $z$-direction. The problem is that both you and your friend used the numbers ${1,2,3,4}$ to label spin sites, which are localized in spacetime. However, the spin sites in each description are non-locally related. Each model assumes a different notion of the system, interactions, and spatial distribution of the physical subsystems. But surely, one model must be the real one, or so you hope.
As you both start to lose patience and commit to discovering what the box hides, you decide to follow the example of particle physicists and grab a hammer to smash it open. Just as you are about to strike the first blow, the box disappears. Nevertheless, the question remains: How can you make sense of what was inside the box?
@@ 2 How space hides in quantum mechanics
The gedanken experiment above illustrates that quantum systems only described by their global Hilbert space, $\mathcal{H}$, a state vector, |$\ket{\Psi}$, and a Hamiltonian, $H$, might not fully characterize its classical properties, such as the notion of a “site” – a location in space [1]. After all, we only had access to the overall energy associated with each of the global states of the system [^4] Once again, we considered a simplified scenario where each energy value displayed in the spectrometer is associated with a non-degenerate energy eigenstate, thus also fixing the size of the Hilbert space. All the eigenvalues together can be used to reconstruct the Hamiltonian matrix in its diagonal form.. This is puzzling because the situation is typically reversed: we impart classical notions of location into our quantum mechanical descriptions. For example, for every spin site located at position $I$, we assign a $\mathbb{C}^{2}$-Hilbert space factor, such that the global Hilbert space is
$\mathcal{H}=\bigotimes_{I=1}\mathbb{C}^{2}\,.$ (2.1)
This is just a simple example of a general attitude [^5] We can speculate whether this attitude is intrinsically derived from the quantization program, where classical observables are promoted to quantum operators, and among them, position and momentum take a special role. Had the program been classicalization instead, then the emphasis could have been rather different.: we typically assign a Hilbert space $\mathcal{H}_{\rm system}$ to each classical subsystem in a given spatial (or more generally, spacetime) region [^6] As usual, this story is fairly more complicated, in particular in the context of infinite-dimensional systems and constrained theories, where only the notion of localizability for the observables is guaranteed [2]. We will come back to that in Sec. 5 and 6. $U$. Thus, some notion of space, or at very least, some notion of location in space, seems to leak into our quantum-mechanical descriptions. Let’s make that explicit.
@@
Thus, we have two commuting sets of observables. If we also require that they together span all the possible operators associated with the system [^7] Remember that any operator can be decomposed into a complex sum of two self-adjoint operators, i.e., observables., i.e., $\mathcal{O}^{A}\otimes\mathcal{O}^{B}\cong\mathcal{L}(\mathcal{H})$, then it follows [5] that the Hilbert space of the quantum system can be factorized, $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. Thus, our local assignment of Hilbert space factors to parts of a bigger system is born out of macrocausality, operationally implemented in a way that leads to microcausality, and finally leading to the factorization of the Hilbert space [^8] Of course, most of the time, we do not think in these terms as we consider non-relativistic quantum mechanics. Then, the factorization of the Hilbert space is simply meant to characterize the independence between quantum subsystems, regardless of a consistent spacetime embedding where state preparations and measurements are causally independent..
Now that the physical intuition is properly encoded in the math, let’s be more precise and general. Following [5], we consider a finite-dimensional Hilbert space $\mathcal{H}$ and a collection of subalgebras of observables corresponding to each subsystem, $\mathcal{A}_{I}$, satisfying the following properties:
  • Local acessibility: Each $\mathcal{A}_{I}$ corresponds to a set of controllable observables;
  • Subsystem independence: $[\mathcal{O}^{I},\mathcal{O}^{J}]=0$ for $I\neq J$, where $\mathcal{O}_{I}\in\mathcal{A}_{I}$;
  • Completeness: $\bigvee_{I}\mathcal{A}^{I}\cong\mathcal{L}(\mathcal{H})$, there is a minimal inclusion of the subalgebras that is isomorphic to the whole algebra of operators.
@@
By choosing a TPS, we determine how the Hilbert space of the entire quantum mechanical system (like the box in the example above) will be partitioned. However, now we can understand that this assignment is intrinsically related to our classical notion of spacetime. On the other hand, if nature is fundamentally quantum mechanical, it might make more sense to start with the global Hilbert space and then find an appropriate decomposition [^9] There are recent works in this direction, for instance [7, 8, 6], while [9] poses for the first time the question of which TPS corresponds to the partition of the Hilbert space into “regions of space”. - particularly one that aligns with our classical observables, such as the notion of site 1 is here, site 2 is there. The question, however, is whether there are multiple classical possibilities associated with the same quantum system, as seen in the different site attributions in our gedanken experiment above, and vice-versa.
Once a TPS is introduced, we examine the 3-tuple $(\mathcal{H},H,\ket{\Psi})$ through the composition of each factor $\mathcal{H}_{I}$ (or local subalgebras, $\mathcal{A}_{I}$). This is exactly what happened in our box experiment, where you and your friend tacitly assumed different TPSs. Each $\mathcal{T}$ considered is not related by local unitaries or permutations of subsystems; instead, the set of variables is non-locally related as shown in (1.3), amounting to a global unitary transformation. Thus, the TPSs are not equivalent. That can also be checked by fixing the factorization of the Hilbert space, $\bigotimes_{I}\mathcal{H}_{I}$ and showing that for each TPSs $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, the operators $T_{1}HT_{1}^{-1}$ and $T_{2}HT_{2}^{-1}$ are not the same up to conjugation by local unitaries, permutations of subsystems, and transposition. Therefore, a change of TPS has an equivalent perspective where the TPS is fixed but the Hamiltonian is unitarily transformed, keeping its spectrum [6].
@@
You might expect that the issue of identifying the appropriate classical limit, understood as the assignment of Hilbert space factors to spatiotemporally localized systems, arises only when dual descriptions exist. After all, if the box cannot be opened (i.e., we cannot extend its Hamiltonian [10]), we have incompatible classical descriptions of what is in the box. However, the problem runs deeper. Even without dualities, a generic Hilbert space with a generic Hamiltonian will have multiple tensor factorizations, only constrained by the dimension of the Hilbert space. And without further specification, all these decompositions are equally reasonable representations of the quantum-mechanical system [1, 11].
This quantum democracy has led to recent efforts to identify criteria for singling out a preferred factorization of the Hilbert space [7, 8]. In particular, [6] suggests a criterion based on the fact that for a TPS with $N$ subsystems, whenever $N\gg k$, the set of $k$-local Hamiltonians has measure zero. If one can then prove that dualities are not present, this would imply an almost unique TPS for a $k$-local Hamiltonian. Thus, by considering $k$-locality of the Hamiltonian as an additional structure to $(\mathcal{H},H,\ket{\Psi})$, a preferred factorization is selected. This also determines a preferred way to map quantum subsystems, as factors in Hilbert space, into classical ones, as systems localized in spacetime [^10] For a critical view of these results, see [12, 1, 11].
@@ 3 A Tale of Emergence
In this section, we will discuss the concept of emergence, start to clarify the distinction between classical and quantum systems, and establish the foundations to better understand our mysterious box.
@@
Key to this definition are the terms novel, robust, and new predictions [^11] On the other hand, one could think that we want to avoid such descriptions by being vague, and seek out definitions of emergence that are strictly mathematical. A recent work in that direction is [16]. We plan to consider this approach in future works.. When we have a description that is novel and robust, in the sense that it stands on its own and provides a way to predict the behavior of physical systems, we typically call it a physical theory. Of course, this also underlies what we mean by a more fundamental description. Therefore, perhaps we should stop here and scrutinize the structure of what we mean by physical theories to make the different flavors of emergence more explicit.
@@ 3.1 The Structure of Physical Theories
  • Underlying state space ($\Omega$): This represents the space of all possible states of a physical system. For example, in Lagrangian mechanics, this is the configuration space, while in Hamiltonian mechanics, it is the phase space. In quantum mechanics, we typically consider the Hilbert space [^12] See footnote 23, though.;
@@
  • Dynamical equation ($U$): Once a state is known for a given degree of freedom, the dynamical equation determines the evolution of the state, defining a trajectory in the state space. In classical mechanics, we have Newton’s, Euler-Lagrange, or Hamiltonian equations [^14] In this section, we will assume classical mechanics is governed by Newton’s equations, as this is sufficient to make our point., while in quantum mechanics, we have the Schrödinger or Heisenberg equation.
In the language of [17], these elements are reduced to a triple $T=(S,Q,D)$: the state-space, the quantities, and the dynamics. The reason why we say most physical theories is that some do not fit this decomposition: e.g., in thermodynamics, the second law does not take the form of a dynamical equation. Nonetheless, to convey our points in the next sections, we will ignore these subtleties.
@@
Now that we have at hand the definition of emergence and the underlying structure of physical theories, we can establish different levels of emergence by considering some examples.
@@ 3.2 Emergence - Level 0: Classical Mechanics and the Center of Mass
$X_{\rm CM}:\left\{\begin{aligned} &\mathds{R}^{3N}\longrightarrow\mathds{R}^{3}\\ &\vec{x}_{\rm CM}\equiv\frac{1}{M}\sum_{I=1}^{N}m_{\rm I}\vec{x}_{\rm I}\,,\quad\quad M=\sum^{N}_{I=1}m_{\rm I}\,,\end{aligned}\right.$
@@
Dynamical Equations $\vec{F}_{I}(\vec{x}_{\rm I})=m_{\rm I}\frac{d\vec{v}_{\rm I}}{dt}$ $\vec{F}(\vec{x}_{\rm CM})=M\frac{d\vec{v}_{\rm CM}}{dt}$
@@ 3.2.1   Does the center of mass exist?
As we aim to build our intuition around ontology as emergence unfolds, we can begin by asking some fundamental questions. There is plenty of literature around this topic [18, e.g. and references therein], but we just want to highlight a few things to slowly build our intuition for later on.
First, in some situations, the center of mass defines a point that lies within the composite object it represents, making it materially instantiated. In such cases, there is literally a material point with its own position and velocity vector that follows the center-of-mass trajectory. However, this is not always the case. Often, the center of mass does not correspond to any actual material point within the defining object. Nonetheless, from a substantivalist perspective, all points in space exist in an absolute sense in any case, so the center of mass trajectory simply labels some of these points.
Rather than focusing on material instantiation or substantivalism to argue for the existence of the center of mass, we can adopt the perspective supported by our definition of emergence. As long as new entities are novel and stable enough to yield reliable predictions, as is the case for the center of mass, we can attribute emergent ontological status to them, similar to Wallace’s implementation [18] of Dennett’s real patterns [19] in physics.